OpenStudy (anonymous):

0.08x + 0.11(500-x) = 50.0

OpenStudy (anonymous):

HELP!

OpenStudy (mathstudent55):

0.08x + 0.11(500-x) = 50.0 Distribute the 0.11 0.08x + 55 - 0.11x = 50 Combine like terms -0.03x + 55 = 50 Subtract 55 from both sides -0.03x = -5 Divide both sides by -0.03 x = -5/(-0.03) x = 166.666... x = 166 2/3 x = 500/3

OpenStudy (anonymous):

thank y ou!

OpenStudy (mathstudent55):

wlcm